# Optical Thermodynamics Beyond the Weak Nonlinearity Limit

### v2

@@ Abstract
Introduction – Multimoded nonlinear photonic systems have been a fertile testbed for the study of many fundamental phenomena, as well as a principal component of many modern technologies. Examples pertaining to the former include foundational questions on wave-turbulence and solitons, optical, BEC and many-body interactions, while on the technological side, one can refer to high-speed communication systems and high-power fiber-based light sources to imaging and cavity microcombs for precision-metrology [5, 7, 27, 13, 2, 11, 3, 25, 38, 34, 18, 22, 8, 33, 29, 33, 32, 17, 41, 26, 1, 42, 45, 39, 40]. In many such frameworks, the intricate nonlinear mode-mode interactions play a crucial role in dictating the light propagation. Though brute-force computations can be invoked for the description of light propagation, they rapidly become power-expensive as the number of modes increases and fail to provide an illuminating description of the underlying beam dynamics. Fortunately, optical thermodynamics (OT) [43, 30, 19, 4, 37, 28, 23, 9] has recently been developed as a new physical framework that circumvents the costly computations while self-consistently describing the complex processes of energy and power exchange among the modes.
@@
At the same time, traditional statistic mechanics has demonstrated that when strong particle-particle interactions are taken into account, new physical phenomena emerge. Examples include the van der Waals and virial corrections to the equation of state, phase transitions, screening, and many-body renormalization, etc [15, 24].
@@
Modeling of nonlinear multimoded photonic circuits – The dynamics of nonlinear multimode photonic networks can be modeled using the framework of time-dependent coupled mode theory. The prototypical associated discrete nonlinear Schrödinger (DNLS) model encompassing diffraction and the cubic nonlinearity of the Kerr effect [16, 10]
@@
where NN is the number of nodes (sites) of the network, and $\psi_{k}$ represents the complex field at each site $k$. Here, we consider only nearest-neighbor couplings with $J=1$, and a Kerr nonlinearity with strength $\chi$. Equation (1) describes coupled single-mode resonators, coupled single-mode waveguide arrays, or multicore fibers [16]. In the former case, the overdot represents a derivative with respect to time $t$ while in the latter two settings it can be thought of as a derivative with respect to the paraxial propagation distance $z$.
@@
Although Eq. (13) has been derived in the strong nonlinear limit $z\rightarrow 0$, remarkably, our computations suggest that the scaling Eq. (14) applies for the whole range of $z$. In fact, the functional form Eq. (13) turns out to be valid in all regimes (including the crossover), thus providing a compact summary of the crossover of the compressibility factor from the weak to the strong nonlinear regime. Our numerical results for various $\beta,\mu,\chi$ values are reported in Fig. 2 and, accordingly, confirm the scaling ansatz Eq. (14) and the applicability of the functional form given by Eq. (13).

### v3

@@ Abstract
Introduction – Multimoded nonlinear photonic systems have been a fertile testbed for the study of many fundamental phenomena, as well as a principal component of many modern technologies. Examples pertaining to the former include foundational questions on wave-turbulence and solitons, optical, BEC and many-body interactions, while on the technological side, one can refer to high-speed communication systems and high-power fiber-based light sources to imaging and cavity microcombs for precision-metrology [6, 8, 30, 16, 2, 14, 3, 28, 42, 38, 21, 25, 9, 36, 32, 36, 35, 20, 45, 29, 1, 46, 49, 43, 44]. In many such frameworks, the intricate nonlinear mode-mode interactions play a crucial role in dictating the light propagation. Though brute-force computations can be invoked for the description of light propagation, they rapidly become power-expensive as the number of modes increases and fail to provide an illuminating description of the underlying beam dynamics. Fortunately, optical thermodynamics (OT) [47, 33, 22, 4, 41, 31, 26, 12, 50, 5] has recently been developed as a new physical framework that circumvents the costly computations while self-consistently describing the complex processes of energy and power exchange among the modes.
@@
At the same time, traditional statistical mechanics has demonstrated that when strong particle-particle interactions are taken into account, new physical phenomena emerge. Examples include the van der Waals and virial corrections to the equation of state, phase transitions, screening, and many-body renormalization, etc [18, 27].
@@
Modeling of nonlinear multimoded photonic circuits – The dynamics of nonlinear multimode photonic networks can be modeled using the framework of time-dependent coupled mode theory. The prototypical associated discrete nonlinear Schrödinger (DNLS) model encompassing diffraction and the cubic nonlinearity of the Kerr effect [19, 13, 10, 37]
@@
where $N$ is the number of nodes (sites) of the network, and $\psi_{k}$ represents the complex field at each site $k$. Here, we consider only nearest-neighbor couplings with $J=1$, and a Kerr nonlinearity with strength $\chi$. Equation (1) describes coupled single-mode resonators, coupled single-mode waveguide arrays, or multicore fibers [19]. In the former case, the overdot represents a derivative with respect to time $t$ while in the latter two settings it can be thought of as a derivative with respect to the paraxial propagation distance $z$.
@@
Although Eq. (13) has been derived in the strong nonlinear limit $z\rightarrow 0$, remarkably, our computations suggest that the scaling Eq. (14) applies for the whole range of $z$. In fact, the functional form Eq. (13) turns out to be valid in all regimes (including the crossover), thus providing a compact summary of the crossover of the compressibility factor from the weak to the strong nonlinear regime. Our numerical results for various $\beta,\mu,\chi$ values are reported in Fig. 2 with the corresponding data available at [11] and, accordingly, confirm the scaling ansatz Eq. (14) and the applicability of the functional form given by Eq. (13).